OpenStudy (anonymous):

Explain how the Quotient of Powers was used to simplify this expression. 5 to the fourth power, over 25 = 52

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

$$\Huge\frac{5^4}{25}=\frac{5^4}{5^2}$$